# Quiver moduli and quantum loop algebras

### v1

@@ 2.4. Cell decompositions
It is well-known that the existence of the cell decomposition (39) implies that the 0-th topological $K$-theory (with $\mathbb{Q}$}-coefficients, in this Section) of $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$ satisfies
@@ 2.7. Interlude on quiver Grassmannians
$\emph{Gr}_{{\boldsymbol{n}}}(T)=\Big\{\left(T_{i}\stackrel{{\scriptstyle\pi_{i}}}{{\twoheadrightarrow}}{\mathbb{C}}^{n_{i}}\right)_{i\in I}\text{ compatibly with the }\phi_{e}\text{'s}\Big\}\Big/\prod_{i\in I}GL_{n_{i}}$
where the compatibility condition means that every $\phi_{e}$ takes $\emph{Ker }\pi_{i}$ to $\emph{Ker }\pi_{j}$ (in other words, the quotients induce an $\boldsymbol{n}$}-dimensional quiver representation on $\{{\mathbb{C}}^{n_{i}}\}_{i\in I}$).
@@
$\emph{Gr}_{{\boldsymbol{n}}}\left(\bigoplus_{i\in I}(I_{i}^{*})^{\oplus m_{i}}\right)\cong M^{\emph{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$
@@
Even though $I_{i}^{*}$ is infinite-dimensional in general, we note that its quiver Grassmannians are finite-dimensional. This is because $I_{i}^{*}$ is generated by the action of ${\mathbb{C}}Q$ on the empty path $\varnothing_{i}$ at $i$, and so a map of quiver representations
$I_{i}^{*}\rightarrow({\mathbb{C}}^{n_{j}})_{j\in I}$
is determined by the vector in ${\mathbb{C}}^{n_{i}}$ where $\varnothing_{i}$ is sent, and by the induced quiver representation on $\{{\mathbb{C}}^{n_{j}}\}_{j\in I}$. More generally, a map of quiver representations
(43) $\bigoplus_{i\in I}(I_{i}^{*})^{\oplus m_{i}}\xrightarrow{\pi}({\mathbb{C}}^{n_{j}})_{j\in I}$
is determined by $m_{i}$ vectors in ${\mathbb{C}}^{n_{i}}$ for all $i\in I$, and by the induced quiver representation on $\{{\mathbb{C}}^{n_{j}}\}_{j\in I}$. This precisely corresponds to a framed quiver representation, and the surjectivity of $\pi$ is equivalent to stability.
@@
$\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}\left(\bigoplus_{i\in I}(I_{i}^{*})^{\oplus m_{i}}\right)\cong M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab},\text{nilp}}$
@@
(44) $\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}\left(\bigoplus_{i\in I}(I_{i}^{*})^{\oplus m_{i}}\right)$ $=\Big\{{\boldsymbol{n}}\text{-dim nilpotent subreps of }\bigoplus_{i\in I}(I_{i}^{**})^{\oplus m_{i}}\Big\}$
$=\Big\{{\boldsymbol{n}}\text{-dim subreps of }\bigoplus_{i\in I}I_{i}^{\oplus m_{i}}\Big\}$
@@ 2.10. Residues
where ${\mathcal{L}}_{1},\dots,{\mathcal{L}}_{n-1}$ are the line bundles on $M^{\text{stab}}_{i_{1},\dots,i_{n-1},{\boldsymbol{m}}}$ that parameterize the lines spanned by the vectors $v_{1},\dots,v_{n-1}$, respectively. The aforementioned claim is straightforward, and left to the reader.
@@ 3.1. Basic definitions
where $d_{ii}$ are positive integers with greatest common divisor 2, and $d_{ij}=d_{ji}$ are non-positive integers $\forall i\neq j$. This data corresponds to a Kac-Moody Lie algebra $\mathfrak{g}$}, which we will refer to as the “type” of all algebras considered hereafter. $I$ is interpreted as a set of simple roots of $\mathfrak{g}$}, hence $\mathbb{Z}^{I}$} is interpreted as the root lattice. The Kac-Moody Lie algebra $\mathfrak{g}$} is called finite type if the symmetric matrix $\left(d_{ij}\right)_{i,j\in I}$ is positive definite, since in this case $\mathfrak{g}$} is a finite-dimensional semisimple Lie algebra. Similarly, $\mathfrak{g}$} is called simply-laced if the generalized Cartan matrix (60) satisfies
(61) $d_{ii}=2,\quad\ \forall i\in I$
@@ 3.2. The big shuffle algebra
whose domain is the negative half (in Drinfeld’s new realization, [10]) of the quantum loop algebra $U_{q}(L{\mathfrak{g}})$ of $\mathfrak{g}$}. We refer to [37] for an overview of the generators-and-relations presentation of $U_{q}(L{\mathfrak{g}})$, and note that it involves sufficiently many relations so that $\Upsilon$ is injective. The corresponding set of relations is so far only known for finite type ([2, 10]) and simply-laced ([35]) $\mathfrak{g}$}, see [36] for a discussion.
@@ 3.3. The (small) shuffle algebra
The inclusion $\subseteq$ of (68) was established in [13] based on the seminal work [14], while the inclusion $\supseteq$ of (68) was proved in [39].
@@ 3.6. Simple modules and $q$-characters
with various $l_{i,\gamma},m_{i,\gamma}\geq 0$ (almost all of which are actually equal to 0) one considers the corresponding simple module $L({\boldsymbol{\psi}})$ of the Borel half of $U_{q}(L{\mathfrak{g}})$ (defined in [24] in the finite type case and in [37] in the arbitrary Kac-Moody case). It was shown in [38, Proposition 4.7] that there exists an isomorphism of vector spaces
(74) $L({\boldsymbol{\psi}})\cong L^{\neq 0}({\boldsymbol{\psi}})\otimes L^{\textbf{ord }{\boldsymbol{\psi}}}$
The second tensor factor in the RHS above was calculated in [38] for finite type $\mathfrak{g}$} (proving a conjecture of [30, 47]) and we will not be concerned with it. Instead, we will focus on the first tensor factor in the RHS of (74): it was realized in [38, Definition 4.3, Remark 4.5] as
@@ 3.7. A model for shuffle algebras
As explained in Subsection 1.2, we are interested in finding a geometric incarnation for the numbers
@@
does not vanish at $(z_{1},\dots,z_{n})=(x_{1},\dots,x_{n})$, and thus the principal ideal $(p)$ together with $((z_{1}-x_{1})^{N},\dots,(z_{n}-x_{n})^{N})$ generate the unit ideal.
@@ 4. Relation to critical $K$-theory
as $\boldsymbol{n}$} ranges over $\mathbb{N}^{I}$} and $\boldsymbol{x}$ ranges over $({\mathbb{K}}^{*})^{{\boldsymbol{n}}}$. We will now relate the vector spaces (86) to critical $K$-theory, which we will argue provides a suitable generalization of Conjecture 1.4 beyond reachable simple modules. We assume that $\mathfrak{g}$} is of finite type throughout the present Section.
@@ 4.1. Quiver with potential
    ([40, Proposition 3.6]) whose image is contained in the image of the push-forward
@@
    obtained by putting together all the maps (90) is an algebra homomorphism, where $\mathcal{V}$} is endowed with the shuffle product (63). Note that this requires appropriately twisting the convolution product, so as to match our specific convention of the zeta functions in (64) on the nose, instead of merely up to a monomial in $x$.
@@
If $\mathfrak{g}$} is of finite type, then $\emph{Im }\iota={\mathcal{S}}$.
@@
where the left-most $({\mathbb{C}}^{*})^{2-c_{ij}}$ parameterizes the arrow maps in (95) and the right-most $({\mathbb{C}}^{*})^{2-c_{ij}}$ acts on it by conjugaction. Restricting a Laurent polynomial in ${\mathcal{V}}_{{\boldsymbol{n}}}\cong K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})$ to $U$ has the effect of performing the specialization (69).
@@ 4.7. Conjectures
Lemma 4.8 implies that we may realize the vector space (23) as the image of the composition $\downarrow\downarrow\rightarrow\rightarrow$ in (106), and we will argue that this composition is more natural than $\rightarrow\rightarrow\downarrow\downarrow$ from a geometric point of view. Firstly, we conjecture that
(107) $\pi\text{ is surjective}$
because it is a composition of two maps: the first is hoped to be surjective due to it being restriction to the stable locus, while the second is certainly surjective due to equivariant localization. Secondly, we conjecture that
(108) $\iota_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{gr},\text{stab}}\text{ is injective}$
since such statements have been proved in other cases (e.g. [7, Theorem 10.2] in cohomology). Once one establishes (107) and (108), then Theorem 1.8 and Lemma 4.8 imply that we have an isomorphism of vector spaces
(109) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\cong\Big[\Delta_{{\mathbf{l}}}^{\text{gr}}\cdot K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\Big]\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
While contingent on (107) and (108), formula (109) provides the natural generalization of the Hernandez-Leclerc conjecture 1.4: since critical $K$-theory of quivers with potential is a natural replacement of the $K$-theory of the critical locus, then
$K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\quad\textbf{is a replacement of}\quad K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}}\cap\text{Crit}(\text{Tr}(W^{\text{gr}})))$
Meanwhile, the multiplication by $\Delta_{{\mathbf{l}}}^{\text{gr}}$ in (109) is simply a derived version of imposing the vanishing of $l_{i,s}$ generic linear combinations of paths ending at any vertex $(i,s)$. Therefore, the RHS of (109) is a natural replacement of $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}$.

### v2

@@ 2.4. Cell decompositions
It is well-known that the existence of the cell decomposition (39) implies that the 0-th $K$-theory (with $\mathbb{Q}$}-coefficients, throughout this Section) of $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$ satisfies
@@ 2.7. Interlude on quiver Grassmannians
(43) $\emph{Gr}_{{\boldsymbol{n}}}(T)=\Big\{\left(T_{i}\stackrel{{\scriptstyle\pi_{i}}}{{\twoheadrightarrow}}{\mathbb{C}}^{n_{i}}\right)_{i\in I}\text{ compatibly with the }\phi_{e}\text{'s}\Big\}\Big/\prod_{i\in I}GL_{n_{i}}$
where the compatibility condition means that every $\phi_{e}$ takes $\emph{Ker }\pi_{i}$ to $\emph{Ker }\pi_{j}$ (in other words, the quotients induce a quiver representation structure on $\{{\mathbb{C}}^{n_{i}}\}_{i\in I}$).
@@
(44) $\emph{Gr}_{{\boldsymbol{n}}}\left(\bigoplus_{k\in I}(I_{k}^{*})^{\oplus m_{k}}\right)\cong M^{\emph{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$
@@
Recall that $I_{k}^{*}$ is freely generated by the action of ${\mathbb{C}}Q$ on the empty path $\varnothing_{k}$ at $k$, hence $(I_{k}^{*})_{i}$ has a basis indexed by paths $p:k\rightarrow i$. Thus, to give surjections
(45) $\left(\bigoplus_{k\in I}\underbrace{(I_{k}^{*})_{i}\oplus\dots\oplus(I_{k}^{*})_{i}}_{m_{k}\text{ summands}}\twoheadrightarrow{\mathbb{C}}^{n_{i}}\right)_{i\in I}$
one needs to provide spanning sets
(46) $\Big(v_{p}^{(k,1)},\dots,v_{p}^{(k,m_{k})}\Big)_{\text{paths }p:k\rightarrow i}\quad\text{of}\quad{\mathbb{C}}^{n_{i}}$
for all $i\in I$. In order for this assignment to be compatible in the sense of (43), the induced quiver maps $\phi_{e}:{\mathbb{C}}^{n_{\text{head}(e)}}\rightarrow{\mathbb{C}}^{n_{\text{tail}(e)}}$ must satisfy
$\phi_{e}(v_{p}^{(k,a)})=\begin{cases}v_{ep}^{(k,a)}&\text{if }\text{tail}(p)=\text{head}(e)\\ 0&\text{otherwise}\end{cases}$
for all $k\in I$ and $a\in\{1,\dots,m_{k}\}$. However, the equation above means that the datum (46) is completely determined by the collection of vectors
$\Big(v_{\varnothing_{k}}^{(k,1)},\dots,v_{\varnothing_{k}}^{(k,m_{k})}\Big)\subset{\mathbb{C}}^{n_{k}}$
as $k\in I$ varies, which altogether generate the quiver representation $\{{\mathbb{C}}^{n_{i}}\}_{i\in I}$. In turn, this datum entails the same information as a point of $M^{\text{stab}}_{{\boldsymbol{n}},{\boldsymbol{m}}}$, with $v_{\varnothing_{k}}^{(k,a)}$ being the image of the $a$-th standard basis vector under the framing map ${\mathbb{C}}^{m_{k}}\rightarrow{\mathbb{C}}^{n_{k}}$.
@@
$\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}\left(\bigoplus_{k\in I}(I_{k}^{*})^{\oplus m_{k}}\right)\cong M_{{\boldsymbol{n}},{\boldsymbol{m}}}^{\text{stab},\text{nilp}}$
@@
(47) $\text{Gr}^{\text{nilp}}_{{\boldsymbol{n}}}\left(\bigoplus_{k\in I}(I_{k}^{*})^{\oplus m_{k}}\right)$ $=\Big\{{\boldsymbol{n}}\text{-dim nilpotent subreps of }\bigoplus_{k\in I}(I_{k}^{**})^{\oplus m_{k}}\Big\}$
$=\Big\{{\boldsymbol{n}}\text{-dim subreps of }\bigoplus_{k\in I}I_{k}^{\oplus m_{k}}\Big\}$
@@ 2.10. Residues
where ${\mathcal{L}}_{1},\dots,{\mathcal{L}}_{n-1}$ are the line bundles on $M^{\text{stab}}_{i_{1},\dots,i_{n-1},{\boldsymbol{m}}}$ that parameterize the lines spanned by the vectors $v_{1},\dots,v_{n-1}$, respectively. The aforementioned claim is straightforward, and left to the reader. ∎
@@ 3.1. Basic definitions
where $d_{ii}=2d_{i}$ are positive integers with greatest common divisor 2, and $d_{ij}=d_{ji}$ are non-positive integers $\forall i\neq j$. This data corresponds to a symmetrizable Kac-Moody Lie algebra $\mathfrak{g}$}, which we will refer to as the “type” of all algebras considered hereafter. We interpret $I$ as a set of simple roots of $\mathfrak{g}$}, hence we interpret $\mathbb{Z}^{I}$} as the root lattice. We call $\mathfrak{g}$} finite type if the symmetric matrix $\left(d_{ij}\right)_{i,j\in I}$ is positive definite, since in this case $\mathfrak{g}$} is a finite-dimensional semisimple Lie algebra. Similarly, $\mathfrak{g}$} is called simply-laced if the generalized Cartan matrix (63) satisfies
(64) $d_{i}=1,\quad\ \forall i\in I$
Following the suggestion of David Hernandez, we call $\mathfrak{g}$} strongly symmetrizable if
(65) $d_{ij}\in\{0,-\max(d_{i},d_{j})\}$
for all $i\neq j$. All finite and affine type $\mathfrak{g}$} are strongly symmetrizable, except for $\widehat{{\mathfrak{sl}}}_{2}$.
@@ 3.2. The big shuffle algebra
whose domain is the negative half (in Drinfeld’s new realization, [11]) of the quantum loop algebra $U_{q}(L{\mathfrak{g}})$ of $\mathfrak{g}$}. We refer to [38] for an overview of the generators-and-relations presentation of $U_{q}(L{\mathfrak{g}})$, and note that (by definition) it involves sufficiently many relations so that $\Upsilon$ is injective. The corresponding set of relations is so far only known explicitly for finite type ([2, 11]), more generally for strongly symmetrizable ([40]), and for simply-laced ([36]) symmetrizable Kac-Moody Lie algebras $\mathfrak{g}$}. See [37] for a discussion.
@@ 3.3. The (small) shuffle algebra
The inclusion $\subseteq$ of (72) was established in [14] based on the seminal work [15], while the inclusion $\supseteq$ of (72) was proved in [41]. More generally, the formulas above hold as stated for strongly symmetrizable $\mathfrak{g}$}, as shown in [40, Section 3.25].
@@ 3.6. Simple modules and $q$-characters
with various $l_{i,\gamma},m_{i,\gamma}\geq 0$ (almost all of which are actually equal to 0) one considers the corresponding simple module $L({\boldsymbol{\psi}})$ of the Borel half of $U_{q}(L{\mathfrak{g}})$ (defined in [25] in the finite type case and in [38] in the arbitrary symmetrizable Kac-Moody case). It was shown in [39, Proposition 4.7] that there is a isomorphism of vector spaces
(78) $L({\boldsymbol{\psi}})\cong L^{\textbf{ord }{\boldsymbol{\psi}}}\otimes L^{\neq 0}({\boldsymbol{\psi}})$
The first tensor factor in the RHS above was calculated in [39] for finite type $\mathfrak{g}$} (proving a conjecture of [31, 49]) and we will not be concerned with it. Instead, we will focus our attention on the second tensor factor in the RHS of (78): this factor was realized in [39, Definition 4.3, Remark 4.5] as
@@ 3.7. A model for shuffle algebras
As explained in Subsection 1.2, we are interested in finding a geometric incarnation for the non-negative integers
@@
does not vanish at $(z_{1},\dots,z_{n})=(x_{1},\dots,x_{n})$, and thus the principal ideal $(p)$ together with $((z_{1}-x_{1})^{N},\dots,(z_{n}-x_{n})^{N})$ generate the unit ideal. ∎
@@ 4. Relation to critical $K$-theory
as $\boldsymbol{n}$} ranges over $\mathbb{N}^{I}$} and $\boldsymbol{x}$ ranges over $({\mathbb{K}}^{*})^{{\boldsymbol{n}}}$. We will now relate the vector spaces (90) to critical $K$-theory of quiver varieties, which we will argue provides a suitable generalization of Theorem 1.4 beyond reachable simple modules. We assume that $\mathfrak{g}$} is strongly symmetrizable throughout the present Section; recall that this includes all finite and affine type symmetrizable Kac-Moody Lie algebras except for $\widehat{{\mathfrak{sl}}}_{2}$.
@@ 4.1. Quiver with potential
    ([42, Proposition 3.6]) whose image is contained in the image of the push-forward
@@
    obtained by putting together all the maps (94) is an algebra homomorphism, where $\mathcal{V}$} is made into an algebra with respect to the shuffle product (67).
@@
If $\mathfrak{g}$} is strongly symmetrizable, then $\emph{Im }\iota={\mathcal{S}}$.
@@
where the left-most $({\mathbb{C}}^{*})^{2-c_{ij}}$ parameterizes the arrow maps in (99) and the right-most $({\mathbb{C}}^{*})^{2-c_{ij}}$ acts on it by conjugation. Restricting a Laurent polynomial in ${\mathcal{V}}_{{\boldsymbol{n}}}\cong K^{{\mathbb{C}}^{*}}(M_{{\boldsymbol{n}}})$ to $U$ has the effect of performing the specialization (73).
@@ 4.7. Coda
Lemma 4.8 implies that we may realize the vector space (23) as the image of the composition $\downarrow\downarrow\rightarrow\rightarrow$ in (110), and we will argue that this composition is more natural than $\rightarrow\rightarrow\downarrow\downarrow$ from a geometric point of view. Indeed, it was proved in [5] that
(111) $\pi\text{ is surjective and }\iota_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{gr},\text{stab}}\text{ is injective}$
using their theory of critical stable envelopes. Therefore, the image of the composition $\downarrow\downarrow\rightarrow\rightarrow$ in (110) is isomorphic to the image of the map $\cdot\Delta_{{\mathbf{l}}}^{\text{gr}}$ therein. Thus
which is the content of Corollary 1.10. Since the left-hand side of the above equation categorifies the coefficients of the $q$-character, (112) is precisely the categorification and generalization to all $\boldsymbol{\psi}$} of Theorem 1.4: since critical $K$-theory of quivers with potential is a natural replacement of the $K$-theory of the critical locus, then
$K(Q^{\text{gr}},W^{\text{gr}})^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}\quad\text{is a replacement of}\quad K(M_{\boldsymbol{x},{\boldsymbol{m}}}^{\text{stab}}\cap\text{Crit}(\text{Tr}(W^{\text{gr}})))$
Meanwhile, multiplication by $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is a derived version of imposing the vanishing of $l_{i,s}$ generic linear combinations of paths ending at any vertex $(i,s)$. Therefore, the RHS of (112) is a natural replacement of $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}$.